# DetailMaster: Can Your Text-to-Image Model Handle Long Prompts?

### v1

@@ 3.1.3 Prompt refinement & enhancement
Prompt refinement and filtering.
Following the attribute extraction, we implement a prompt refinement pipeline to transform the original captions into detail-rich long prompts. For each sample, we enumerate all identified main characters and employ an LLM to incorporate both “Character Attributes” and “Structured Character Locations” of each character into the original caption. Subsequently, the same LLM integrates the sample’s “Scene Attributes” and “Explicit Spatial/Interactive Relationships” to produce the refined prompt. Beyond prompt refinement, the extracted attributes are ultimately used to generate evaluation metrics. To guarantee the accuracy of evaluation metric, we implement an attribute validation mechanism where an LLM verifies the presence of each attribute in the refined prompt, filtering out any attributes that fail to be incorporated. In addition, samples with less than four valid character-level attributes are filtered out.
Prompt enhancement.
Throughout the attribute extraction and prompt refinement pipeline, we employ Qwen2.5-14B-Instruct qwen2.5 as our primary LLM and Qwen2.5-VL-7B-Instruct qwen2.5-VL as the MLLM. This initial process yields 4,565 detail-rich prompts, with an average token length of 237.53. To achieve higher precision and greater detail, we subsequently reprocess the corresponding 4,565 samples using Qwen2.5-VL-72B-Instruct as both the LLM and the MLLM. This new process generates an improved dataset containing 4,116 prompts with an average token length of 284.89.
@@ 3.2.1 Overview
Quantitatively measuring the quality of compositional text-to-image generation proves inherently difficult, as it demands fine-grained cross-modal understanding between textual prompts and generated images. Popular approaches typically employ object detection models for spatial verification, vision-language models for image-text alignment clipscore , and MLLMs for attribute verification dsg ; t2i_compbench . However, these methods often yield coarse-grained evaluations that may introduce noise and compromise measurement accuracy. To address these limitations, we develop a robust multi-stage evaluation pipeline that systematically assesses the accuracy of all four categories of attributes.
@@ 3.2.2 Evaluation metrics
Regarding the design of evaluation metrics, we utilize all four categories of attributes that we previously extract. For the “Character Attributes” metric, we count the number of correctly generated attributes for each character, and compute accuracy by dividing the total number of correct attributes by the total attributes of characters per category (categorized as object, animal, or person). The “Character Locations” metric involves tallying the number of correctly positioned characters, with accuracy derived from the ratio of accurately placed characters to all locatable characters. For “Entity Relationships”, we examine each relationship in every sample and calculate the proportion of correct relationships. These three metrics assess attribute mismatches and omissions in detail-rich scenarios. For the “Scene Attributes” metric, we compute accuracy for background, lighting, and style, evaluating overall image fidelity in long prompt scenarios.
In the main text, our evaluations are conducted on the full version of DetailMaster, comprising 4,116 prompts. Additionally, we design a mini version of our benchmark to facilitate rapid evaluation for researchers, which consists of 800 detail-rich prompts. The details are in Appendix H.
@@ 3.2.3 Evaluation pipeline
We design a multi-step evaluation pipeline to systematically assess the performance of text-to-image models. First, we need to instruct the T2I model with the prompts from DetailMaster to generate corresponding image sets. Next, we detect and localize the characters present in the generated images. Subsequently, we conduct a rigorous quantitative evaluation across four critical dimensions: Character Attributes, Character Locations, Scene Attributes, and Entity Relationships, deriving accuracy rates for each evaluated model. To ensure computational efficiency and resource optimization, we employ Qwen2.5-VL-7B-Instruct as the MLLM throughout the evaluation process. Further details of the evaluation pipeline are provided in Appendix E.
@@ 4.1 Experimental setup
We evaluate the performance of seven general-purpose models and five long-prompt optimized models. Among the general-purpose models, five are open-source implementations including Stable Diffusion 1.5 stablediffusion , SD-XL sdxl , DeepFloyd IFimagen , Stable Diffusion 3.5 Large sd3 , and FLUX.1-dev flux2024 , while two are proprietary commercial systems, specifically the image generation modules of Gemini 2.0 Flash Gemini_image and GPT-4o gpt4o-image1 . The long-prompt optimized models include LLM4GEN denseprompt , ELLA DPG-Bench , LongAlign longalign , LLM Blueprint Llm_blueprint , and ParaDiffusion paradiffusion . Detailed configurations for all the evaluated models are provided in Appendix B. Regarding the evaluation metrics, we assess each model across four key generation tasks: Character Attributes, Character Locations, Scene Attributes, and Entity Relationships, enabling systematic assessment of compositional text-to-image generation in long-prompt scenarios. The results are presented in Table 2.
Table 2: Evaluation results on the DetailMaster Benchmark. We evaluate model performance across four key tasks: character attributes, character locations, scene attributes, and entity relationships. All values in the table represent accuracy percentages.
@@
SD1.5 stablediffusion - 20.79 27.69 13.89 8.68 22.02 64.52 80.90 5.88
SD-XL sdxl - 24.41 29.54 16.73 10.95 27.52 68.42 68.87 9.83
DeepFloyd IF DeepFloydIF - 31.01 37.47 25.61 14.28 26.26 67.87 86.49 11.95
SD3.5 Large sd3 - 48.56 46.20 32.95 33.62 89.61 90.33 95.69 40.03
FLUX.1-dev flux2024 - 51.47 45.83 34.91 41.57 95.77 97.05 94.81 47.49
Gemini 2.0 Flash Gemini_image - 55.44 47.84 34.23 44.74 96.69 95.90 97.20 50.78
GPT Image-1 gpt4o-image1 - 59.41 48.04 40.40 53.92 97.50 98.85 97.69 63.07
@@
LLM4GEN denseprompt SD1.5 21.75 29.14 17.20 9.22 25.60 65.71 50.11 7.44
LLM Blueprint Llm_blueprint SD1.5 21.41 27.01 13.91 18.44 50.89 77.57 64.24 11.60
ELLA DPG-Bench SD1.5 34.50 35.14 22.01 15.92 47.12 74.28 40.14 16.64
LongAlign longalign SD1.5 32.95 34.60 15.35 14.89 77.03 87.70 72.27 19.17
ParaDiffusion paradiffusion SD-XL 35.25 33.28 22.29 20.32 83.65 92.21 67.45 25.50
@@ 4.2.1 Comparisons across general-purpose text-to-image models
Impact of token length constraints.
As shown in Table 2, models employing CLIP as the text encoder (SD1.5 and SD-XL) show significant limitations when handling long prompts due to token length constraints, consistently underperforming across all four tasks. In contrast, DeepFloyd, which employs T5 as its text encoder, benefits from an extended input capacity and visibly outperforms the former two models. Nevertheless, its performance remains bad, constrained by both the quality of its training data and inherent architectural limitations.
Performance ceiling of SOTA models.
SD3.5 and FLUX, which combine T5 with enhanced training data and superior model architecture, show improvements over previous models. These two models achieve remarkable performance on “Scene Attributes”, attaining accuracy exceeding 90%, indicating robust scene control capability. However, their performance remains constrained on the other three tasks, suggesting that the absence of explicit training on long prompts limits their adaptability. The proprietary models, Gemini 2.0 Flash and GPT Image-1, outperform all aforementioned open-source models. Nevertheless, their accuracy for “Character Attributes”, “Character Locations”, and “Entity Relationships” also plateaus at around 50%, indicating that even SOTA models have considerable room for improvement in long prompt scenarios.
@@ 4.2.2 Comparisons across long-prompt optimized text-to-image models
Dense prompt training matters more than increasing token capacity.
While LLM4GEN uses an adapter to infuse T5 features, its 128-token limit during training fails to address long prompt challenges, leading to slight improvements. ELLA maintains the same constraint but train with dense and complex prompt, yielding more improvements than LLM4GEN. With a similar architecture and training strategy, ParaDiffusion extends the limit to 512 during training, enabling superior performance. Compared to DeepFloyd (with same token capacity but conventional training), ParaDiffusion shows better performance across almost all metrics. These results validate that while expanded token capacity provides necessary infrastructure, dense prompt training yields greater gains.
Decomposition and iteration mitigates long-prompt challenges.
LongAlign decomposes long prompts for separate encoding and trains with long prompts, indirectly increasing the token capacity and achieving notable gains. LLM Blueprint identifies key roles and locations from long prompts, followed by iterative image refinement, yielding notable improvements in “Character Locations”.
Performance bottlenecks in backbone architectures.
Nevertheless, current long-prompt optimized methods predominantly build upon SD1.5 and SD-XL. While these methods show measurable improvements over their baseline counterparts, their overall performance remains unsatisfactory.
@@
4.2.3 Analysis on key generation tasks
Task-specific performance trends.
Consistent performance trends are shown across generation tasks, with stronger models achieving higher scores. Notably, “Character Attributes”, “Character Locations”, and “Entity Relationships” remain challenging, while “Scene Attributes” proves easier. Interestingly, for the “style” attribute, the optimized models underperform baseline models. This may attribute to the predominance of realistic styles in the dataset: while SD1.5 naturally aligns with real-world styles, optimizations may introduce undesirable deviations.
Stronger models, fewer attribute misalignment and loss.
To further analyze model behavior, we calculate attribute accuracy specifically for the detected generated characters (excluding missing ones), revealing how models handle attribute misalignment and omission. As shown in Table 3, advanced or optimized models show superior attribute accuracy for generated characters, indicating their enhanced prompt comprehension and better attribute-character alignment capabilities.
@@ 4.2.3 Analysis on key generation tasks
To better understand the impact of prompt length on text-to-image generation, we analyze the relationship between prompt token length (tokenized using CLIP’s tokenizer) and the accuracy of “Character Attributes”, “Character Locations”, and “Entity Relationships”. We employ five distinct length intervals: below 250, above 400, and three 50-token intervals spanning 250-400. The results, illustrated in Figure 3, show a consistent negative correlation between prompt length and generation accuracy across 10 evaluated models (excluding SD1.5 and SD-XL). This indicates that current text-to-image models indeed struggle to faithfully interpret and execute long prompts, suggesting substantial room for improvement in handling complex prompt scenarios.
Furthermore, we conduct additional analysis on the relationship between prompt length and attribute misalignment, visualizing how attribute accuracy for generated characters decreases with longer prompts. The details are shown in Appendix G.
@@ 5.1 Data statistics
We present a statistical analysis of our DetailMaster Benchmark. The dataset consists of 4,116 prompts, covering 5,165 distinct nouns with an average prompt length of 284.89 tokens. The token length distribution reveals: 285 prompts at 100-200, 2,399 prompts at 200-300, 1,151 prompts at 300-400, and 281 prompts exceed 400. Regarding evaluation metrics: 1) The “Character Attributes” category includes 8,597 valid characters annotated with 37,165 distinct features (comprising 22,728 for “object”, 4,810 for “animal”, and 9,627 for “person”). 2) The “Character Locations” category contains 6,910 character position annotations. 3) For “Scene Attributes”, we annotate 4,104 background descriptions, 4,114 lighting descriptions, and 4,112 style descriptions. 4) The “Entity Relationships” category contains 18,526 explicit spatial/interactive relationship annotations. As illustrated in Table 1, our benchmark surpasses existing benchmarks for compositional or long-prompt-driven image generation through more comprehensive metrics and longer prompts, enabling rigorous assessment of text-to-image models’ capacities in handling complex, long prompts.
@@ 5.2 Human evaluation
To assess the data quality of our DetailMaster Benchmark, we randomly select 50 samples from each evaluation task (400 in total) and employ two expert annotators to evaluate them. The final scores are derived through an averaging process. The human evaluation criteria comprises three primary dimensions: 1) Task Relevance. We evaluate whether the evaluation data properly aligns with the designated evaluation tasks. 2) Source Image Fidelity. We evaluate whether the evaluation data accurately aligns with the original images from the detailed captioning dataset (for hallucination assessment). 3) Prompt Consistency. We evaluate whether the evaluation content is described in the final polished prompts. The results show that 100% of samples meet task relevance requirements, 93.6% maintain visual fidelity with original images, and 97.5% show complete consistency with final polished prompts. These findings validate that the evaluation data of DetailMaster faithfully derives from authentic image-caption pairs, ensuring its validity. Furthermore, the strong alignment between evaluation data and final prompts mitigates accuracy evaluation errors caused by missing attribute descriptions. These results collectively confirm the high quality and robustness of our benchmark. Detailed human evaluation results for specific sub-tasks are provided in Appendix D.
6 Overview of more experiments and analysis
In Appendix A, we examine our spatial partitioning scheme designed for extracting “Character Locations” features and validate its effectiveness. Appendix B details configurations of all evaluated models. Appendix C documents the resource requirement. Appendix D documents the full human evaluation results, further confirming the high quality of DetailMaster. Appendix E describes implementation details of our evaluation pipeline. Appendix F verifies the non-sensitive impact of randomness when using LLMs in our evaluation. Appendix G studies the positive correlation between prompt length and attribute misalignment. Appendix H presents the mini version of DetailMaster along with corresponding results. Appendix I presents detailed discussion of limitations. Appendix J showcases case studies, validating the correctness of our evaluation conclusions.
7 Conclusion
We present DetailMaster, a benchmark for evaluating text-to-image generation in long prompt scenarios, along with a robust and fine-grained evaluation protocol. Our benchmark comprises 4,116 long and detail-rich prompts, with an average token length of 284.89. The evaluation covers four key generation tasks: “Character Attributes”, “Structured Character Locations”, “Multi-Dimensional Scene Attributes”, and “Explicit Spatial/Interactive Relationships”. And we assess 7 general-purpose and 5 long-prompt-optimized T2I models, comparing their architectural designs and optimization strategies, and derive key insights: generation accuracy declines as prompt length increases, with even SOTA models achieving only $\sim$50accuracy, indicating substantial room for improvement. We hope our proposed benchmark and evaluation protocol will drive advancements in text-to-image models for long prompts, enabling better adherence to user instructions in specialized domains.
Appendix

### v2

@@ 3.1.3 Prompt Refinement & Enhancement
Prompt refinement and filtering. Following the attribute extraction, we implement a prompt refinement pipeline to transform the original captions into detail-rich long prompts. For each sample, we enumerate all identified main characters and employ an LLM to incorporate all their attributes and locations into the original caption. Subsequently, the same LLM integrates the sample’s scene attributes and spatial/interactive relationships to produce the refined prompt.
Beyond prompt refinement, the extracted attributes also form the basis of our evaluation metrics. To ensure metric accuracy, we implement an attribute validation mechanism where an LLM verifies the presence of each attribute in the refined prompt, filtering out any unincorporated attributes. In addition, samples with less than 4 valid character-level attributes are filtered out. (see Figure 2 ⑥)
Prompt enhancement. Throughout the attribute extraction and prompt refinement pipeline, we employ Qwen2.5-14B-Instruct (Team, 2024) as our primary LLM and Qwen2.5-VL-7B-Instruct (Team, 2025) as the MLLM. This initial process yields 4,565 detail-rich prompts, with an average token length of 237.53. To achieve higher precision and greater detail, we subsequently reprocess the corresponding 4,565 samples using Qwen2.5-VL-72B-Instruct as both the LLM and the MLLM. This new process generates an improved dataset containing 4,116 prompts with an average token length of 284.89. (see in Figure 2 ⑦)
@@ 3.2.1 Overview
Evaluating compositional T2I generation proves inherently difficult, as it demands fine-grained cross-modal understanding between textual prompts and generated images. Popular approaches typically employ object detection models for spatial verification, vision-language models for image-text alignment (Hessel et al., 2021), and MLLMs for attribute verification (Cho et al., 2023; Huang et al., 2023). However, these methods often yield coarse-grained and noisy assessments, compromising evaluation accuracy. To address these limitations, we develop a robust multi-stage evaluation pipeline that systematically assesses the accuracy of all four categories of attributes.
@@ 3.2.2 Evaluation Metrics
Our evaluation framework is built upon the four categories of attributes that we previously extract. For the “Character Attributes” metric, we calculate it as the ratio of correctly rendered attributes to the total number of specified attributes for each character category (i.e., object, animal, and person). The “Character Locations” metric assesses accuracy by computing the proportion of characters positioned correctly relative to the total number of annotated characters. For “Entity Relationships”, we calculate the percentage of correctly rendered relationships among all those described in the prompt. These three metrics quantify the model’s performance on attribute-level details, specifically targeting mismatches and omissions. For the “Scene Attributes” metric, we compute accuracy for background, lighting, and style, evaluating overall image fidelity in long prompt scenarios.
Our main evaluations are conducted on the full version of our DetailMaster, comprising 4,116 prompts. Additionally, we design a mini version of our benchmark to facilitate rapid evaluation for researchers, which consists of 800 detail-rich prompts. The details are in Appendix L.
@@ 3.2.3 Evaluation Pipeline
We introduce a multi-step evaluation pipeline to systematically assess the compositional generation capabilities of T2I models. First, we instruct the T2I models with prompts from our DetailMaster benchmark to generate corresponding image sets. Next, we detect and localize the characters present in the generated images. Subsequently, we conduct a rigorous quantitative evaluation across our four critical dimensions, deriving accuracy rates for each evaluated model. Further details of the evaluation pipeline are provided in Appendix F.
On the Robustness of an MLLM-based Evaluator. While leveraging a single MLLM family (i.e., Qwen) for both data curation and evaluation could raise concerns about potential self-enhancement bias, we argue this risk is mitigated for two key reasons. First, our data construction is heavily grounded by auxiliary tools like open-set object detectors and guided by original human-annotated captions, breaking a purely end-to-end LLM pipeline. Second, as detailed in our robustness analysis (Appendix I), re-evaluating all models with a distinct MLLM (i.e., InternVL) preserves the relative model rankings and core conclusions. This confirms that DetailMaster measures a general compositional capability rather than an affinity for a specific model’s idiosyncrasies.
@@ 4.1 Experimental Setup
We evaluate the performance of seven general-purpose models and five long-prompt optimized models. Among the general-purpose models, five are open-source implementations including Stable Diffusion 1.5 (Rombach et al., 2022), SD-XL (Podell et al., 2023), DeepFloyd IF(Saharia et al., 2022), Stable Diffusion 3.5 Large (Esser et al., 2024), and FLUX.1-dev (Labs, 2024), while two are proprietary commercial systems, specifically the image generation modules of Gemini 2.0 Flash (Google, 2025) and GPT-4o (OpenAI, 2025). The long-prompt optimized models include LLM4GEN (Liu et al., 2025), ELLA (Hu et al., 2024), LongAlign (Liu et al., 2024), LLM Blueprint (Gani et al., 2023), and ParaDiffusion (Wu et al., 2023a). Detailed configurations for all the evaluated models are provided in Appendix C. Regarding the evaluation metrics, we assess each model across four key generation tasks: Character Attributes, Character Locations, Scene Attributes, and Entity Relationships, enabling systematic assessment of compositional text-to-image generation in long-prompt scenarios. The results are presented in Table 2.
Table 2: Results on the DetailMaster Benchmark. Values represent accuracy percentages.
@@
SD1.5 (Rombach et al., 2022) - 20.79 27.69 13.89 8.68 22.02 64.52 80.90 5.88
SD-XL (Podell et al., 2023) - 24.41 29.54 16.73 10.95 27.52 68.42 68.87 9.83
DeepFloyd IF (at StabilityAI, 2023) - 31.01 37.47 25.61 14.28 26.26 67.87 86.49 11.95
SD3.5 Large (Esser et al., 2024) - 48.56 46.20 32.95 33.62 89.61 90.33 95.69 40.03
FLUX.1-dev (Labs, 2024) - 51.47 45.83 34.91 41.57 95.77 97.05 94.81 47.49
Gemini 2.0 Flash (Google, 2025) - 55.44 47.84 34.23 44.74 96.69 95.90 97.20 50.78
GPT Image-1 (OpenAI, 2025) - 59.41 48.04 40.40 53.92 97.50 98.85 97.69 63.07
@@
LLM4GEN (Liu et al., 2025) SD1.5 21.75 29.14 17.20 9.22 25.60 65.71 50.11 7.44
LLM Blueprint (Gani et al., 2023) SD1.5 21.41 27.01 13.91 18.44 50.89 77.57 64.24 11.60
ELLA (Hu et al., 2024) SD1.5 34.50 35.14 22.01 15.92 47.12 74.28 40.14 16.64
LongAlign (Liu et al., 2024) SD1.5 32.95 34.60 15.35 14.89 77.03 87.70 72.27 19.17
ParaDiffusion (Wu et al., 2023a) SD-XL 35.25 33.28 22.29 20.32 83.65 92.21 67.45 25.50
@@ 4.2.1 Comparisons across General-purpose Text-to-Image Models
Impact of token length constraints. As shown in Table 2, models employing CLIP as the text encoder (SD1.5 and SD-XL) show significant limitations when handling long prompts due to token length constraints, consistently underperforming across all four tasks. In contrast, DeepFloyd IF, which employs T5 as its text encoder, benefits from an extended input capacity and visibly outperforms the former two models. Nevertheless, its performance remains bad, constrained by both the quality of its short training data and inherent architectural limitations.
Performance ceiling of SOTA models. SD3.5 and FLUX, which combine T5 with enhanced training data and superior model architecture, show improvements over previous models. These two models achieve remarkable performance on “Scene Attributes”, attaining accuracy exceeding 90%, indicating robust scene control capability. However, their performance remains constrained on the other three tasks, suggesting that the absence of explicit training on long prompts limits their adaptability. The proprietary models, Gemini 2.0 Flash and GPT Image-1, outperform all aforementioned open-source models. Nevertheless, their accuracy for “Character Attributes”, “Character Locations”, and “Entity Relationships” also plateaus at around 50%, indicating that even SOTA models have considerable room for improvement in long prompt scenarios.
@@ 4.2.2 Comparisons across Long-Prompt Optimized Text-to-Image Models
Long prompt training matters more than increasing token capacity. While LLM4GEN uses an adapter to infuse T5 features, its 128-token limit during training fails to address long prompt challenges, leading to slight improvements. ELLA maintains the same token constraint but train with long and complex prompt, yielding more improvements than LLM4GEN. With a similar architecture and training strategy, ParaDiffusion extends the limit to 512 during training, enabling superior performance. Compared to DeepFloyd IF (with same token capacity but conventional training), ParaDiffusion shows better performance across almost all metrics. These results validate that while expanded token capacity provides necessary infrastructure, long prompt training yields greater gains.
Decomposition and iteration mitigates long-prompt challenges. LongAlign decomposes long prompts for separate encoding and trains with long prompts, indirectly increasing the token capacity and achieving notable gains. LLM Blueprint identifies key roles and locations from long prompts, followed by iterative image refinement, yielding notable improvements in “Character Locations”.
Performance bottlenecks in backbone architectures. Nevertheless, current long-prompt optimized methods predominantly build upon SD1.5 and SD-XL. While these methods show measurable improvements over their baseline counterparts, their overall performance remains unsatisfactory.
4.2.3 Analysis on Key Generation Tasks
Metric difficulty hierarchy. Notably, “Character Attributes”, “Character Locations”, and “Entity Relationships” remain challenging, while “Scene Attributes” proves easier. We quantitatively establish this hierarchy in Appendix M.
Error source analysis. Our analysis of failure cases (detailed in Appendix M) illuminates the model’s primary weaknesses. The predominant error patterns are: 1) conflicts between spatial instructions and the model’s real-world positional priors; 2) difficulty rendering complex descriptions, including detailed character interactions and intricate apparel; and 3) cascading failures, where initial errors in character attributes preclude the correct generation of entity relationships; 4) errors in scene attributes, typically confined to excessive structural complexity or highly specialized styles.
@@ 4.2.3 Analysis on Key Generation Tasks
Style degradation from optimization. Interestingly, for the style attribute, the optimized models underperform. This may stem from our benchmark’s bias towards the realistic styles: while SD1.5 naturally aligns with real-world styles, optimizations may introduce undesirable deviations.
Stronger models mitigate attribute misalignment. We assess attribute accuracy exclusively for successfully generated characters to isolate model performance on attribute misalignment from character omission. Table 3 shows that advanced models achieve superior accuracy, demonstrating enhanced prompt comprehension and more precise attribute-character alignment.
@@
To better understand the impact of prompt length on text-to-image generation, we analyze the relationship between prompt token length (tokenized using CLIP’s tokenizer) and the accuracy of “Character Attributes”, “Character Locations”, and “Entity Relationships”. We employ five distinct length intervals: below 250, above 400, and three 50-token intervals spanning 250-400. The results, illustrated in Figure 3, show a consistent negative correlation between prompt length and generation accuracy across 10 evaluated models (excluding SD1.5 and SD-XL due to their limited context windows). This indicates that current T2I models indeed struggle to faithfully follow long prompts, suggesting substantial room for improvement in handling complex prompt scenarios.
Furthermore, our analysis in Appendix N reveals this is not merely an issue of character omission; even for the present characters, the fidelity of attribute alignment also deteriorates with increasing prompt length, highlighting a core failure in attribute binding under detail-intensive conditions.
@@ 5.1 Data Statistics
Our DetailMaster Benchmark contains 4,116 prompts, covering 5,165 distinct nouns with an average prompt length of 284.89 tokens. The token length distribution reveals: 285 prompts at 100-200, 2,399 prompts at 200-300, 1,151 prompts at 300-400, and 281 prompts exceed 400. Its comprehensive annotations include: 1) “Character Attributes”: 8,597 valid characters annotated with 37,165 distinct features (22,728 for “object”, 4,810 for “animal”, and 9,627 for “person”); 2) “Character Locations”: 6,910 character position annotations; 3) “Scene Attributes”: 4,104 background descriptions, 4,114 lighting descriptions, and 4,112 style descriptions. 4) “Entity Relationships”: 18,526 spatial/interactive relationship annotations. As illustrated in Table 1, our benchmark surpasses existing benchmarks through more comprehensive metrics and longer prompts, offering a more rigorous test for T2I models on long-form instructions.
@@ 5.2 Human Evaluation
To assess the data quality of our DetailMaster Benchmark, we randomly select 50 samples from each evaluation task (400 in total) and employ two expert annotators to evaluate them, with scores being averaged. The evaluation is based on three criteria: 1) Task Relevance: alignment of the prompts and annotations with their designated tasks; 2) Source Image Fidelity: alignment of the prompts and annotations with the source images (for hallucination assessment); 3) Prompt Consistency: whether the annotations are reflected in the final polished prompts. The results show that 100% of samples meet task relevance requirements, 93.6% maintain visual fidelity with source images, and 97.5% show complete consistency with final polished prompts. These findings validate that our DetailMaster Benchmark faithfully derives from authentic image-caption pairs, ensuring its validity. Furthermore, the strong alignment mitigates evaluation errors caused by missing attribute descriptions. These results collectively confirm the high quality and robustness of our benchmark. Detailed human evaluation results for specific sub-tasks are provided in Appendix G.
6 Further Analysis and Insights
Benchmark Validity and Robustness. We confirm the high fidelity of DetailMaster through expert human evaluation, which shows near-perfect alignment between our prompts, annotations, and the source images (Appendix G). The robustness of our LLM-based evaluation is established by consistent model rankings and conclusions across a different evaluator and random seeds (Appendix I, J), ensuring that our findings are reproducible and not an artifact of the evaluator choice.
A Hierarchy of Compositional Failure. Our fine-grained metrics reveal a clear difficulty hierarchy for T2I models. Precise spatial reasoning (Character Locations) is the most challenging task, which due to conflicts with strong real-world priors, followed by rendering complex Person Attributes (Appendix M). In contrast, general Scene Attributes like lighting and style are rendered with much higher fidelity. This hierarchy pinpoints critical areas for future research.
Attribute Binding Fails Under High Detail Loads. The negative correlation between prompt length and accuracy (Figure 3) is not merely due to character omission. Our analysis shows that even for the present characters, attribute accuracy degrades as prompts grow longer (Appendix N).
Fostering Broader Adoption. To promote community adoption, we provide a resource-efficient Mini-Benchmark for rapid evaluation, which preserves the model ranking trends of the full dataset (Appendix L). Furthermore, we demonstrate the framework’s compatibility with external metrics for assessing orthogonal dimensions such as image aesthetics and safety (Appendix K).
7 Discussion and Conclusion
We introduce DetailMaster, the first large-scale benchmark for evaluating T2I models on long, detail-intensive prompts. Our evaluations with 12 models reveal a critical performance gap: even SOTA systems exhibit deficiencies in complex compositional generation, with prompt adherence deteriorating as prompt length grows; the primary drivers of enhanced long-prompt adherence are not larger context windows but methods such as long-prompt training and iterative decomposition. We hope DetailMaster enables more work towards building more precise, controllable, and detail-oriented T2I models, unlocking applications hindered by a lack of compositional fidelity.
Based on our analysis, we hypothesize the failures in long-prompt scenarios are caused by text encoders that “flatten” compositional grammar and diffusion models that permit “attribute leakage.” Direct causal validation of these mechanisms remains an open question, illuminating several avenues for future research: 1) Structure-Aware Text Encoders: Developing encoders that explicitly model a prompt’s syntactic or semantic graph structure, rather than treating it as a flat token sequence, to better preserve compositional meaning. 2) Object-Centric Diffusion Models: Investigating architectures that learn disentangled, object-centric latent representations (perhaps via specialized attention or binding slots) to mitigate attribute leakage and enforce stricter object-attribute associations. 3) Curriculum-Based and Data-Centric Training: Scaling our data generation pipeline to create massive, high-quality datasets for long-prompt training, and exploring curriculum learning that progresses from simple to complex prompts to improve compositional generalization.
Appendix A Overview of the Appendix